OpenStudy (anonymous):

A family is building a new house on a rectangular lot of land. The lot measures 7x by 11x and the house will be 2x by 3x. How much land will be left for their yard after the house is built? A: 77^2 B: 83x^2 C: 40x^2 D: 71x^2

OpenStudy (luigi0210):

Well the total area is \[\LARGE7x*11x=77x^2\] What's being used is \[\LARGE 2x*3x=6x^2\] I would assume that you subtract the areas? \[\LARGE 77x^2-6x^2=71x^2\]

OpenStudy (anonymous):

thanx

OpenStudy (luigi0210):

Yup, and welcome to Openstudy btw :)

OpenStudy (anonymous):

sorry I forgot my password can you help me with somemore or maybe u can add me on fb and I can send you pics and itll help you better